stu144639:

A population of mice has black or brown hair. The gene for black hair color (B) is dominant over brown (b). If 16% of the mice are homozygous black and 24% are heterozygous black, what would be the frequency of each allele? (p + q = 1, p2 + 2pq + q2 = 1)

Vocaloid:

Homozygous Black = b^2 = .16 Solve for b Then use 1 - b = B to find the frequency of the dominant allele